# Free Majorana Fermion Meets Gauged Ising Conformal Field Theory on the Fuzzy Sphere

### v2

@@ Abstract
The fuzzy sphere regularisation is a powerful tool to study conformal field theories (CFT) in three spacetime dimensions. In this paper, we extend its scope to CFTs with local fermionic operators. We realise the free Majorana fermion CFT on a setup with one flavour of bosons and one flavour of fermions on the lowest Landau level with a $1/2$ angular momentum mismatch. We allow conversion between two bosons and two fermions, and use a relative chemical potential as the tuning parameter. On the phase diagram, we find three gapped phases, viz. a fermionic integer quantum Hall phase, an $f$-wave chiral topological superconductor, and a bosonic Pfaffian phase. They are separated by two continuous transitions described respectively by a free Majorana fermion and a gauged Ising CFT. We numerically confirm their emergent conformal symmetry through the operator spectrum. We also study the two-point correlation function of the Majorana fermion, and the topological Wilson line defect of the gauged Ising CFT. Our work lays the foundation for studying interacting fermionic CFTs and regularising quantum field theories on the fuzzy sphere.
@@ 1 Introduction
Conformal field theory (CFT) is one of the central topics of modern physics. CFT has provided important insights into various aspects of theoretical physics. In condensed matter physics, it has produced useful predictions about the critical phenomena [1, 2, 3]. Many classical and quantum phase transitions are conjectured to have emergent conformal symmetry in the infrared (IR). In high energy physics and quantum field theories, CFT also plays important role in understanding string theory [4], duality with gravitational theories on anti-de Sitter (AdS) space [5], and renormalisation group flow structure [6]. In 2D, the infinite-dimensional conformal algebra has made many theories exactly solvable [7, 8, 9]. However, going to the higher dimensions, the CFTs are less well-studied due to a much smaller conformal group. Some of the conformal data are determined at high precision by conformal bootstrap [10, 11].
@@
Fermionic CFTs, i. e. CFTs that include fermionic local operators, constitute a particularly interesting type of CFTs, including Gross-Neveu-Yukawa theories and certain gauge theories with or without Chern-Simons (CS) term. In some special cases, these CFTs can have emergent supersymmetry. For example, the theory of a single Majorana fermion coupled to a real Yukawa scalar field flows to a superconformal field theory with $\mathscr{N}=1$ in the IR, known as the super-Ising theory [12, 13, 14]. Experimentally, these fermionic CFTs are easier to detect than their bosonic counterparts, as they produce rich electric signals, to which most experimental probes are naturally sensitive. For example, the frequency dependence of conductivity can be measured through electric transport experiments, and it corresponds to the OPE of the conserved symmetry current $J\times J$ [15, 16], including the central charge $C_{J}$ from its two-point function. An outstanding platform is the Moiré materials [17, 18, 19]. Through tuning many available knobs like potential amplitude, they can be used to realise plateau transitions between fractional Chern insulators [20, 21, 22]. However, the nature and critical data of these theories remain poorly understood.
@@
The fuzzy sphere regularisation [23] has emerged as a new powerful method to study 3D CFTs. This method involves studying quantum systems on a sphere that is ‘fuzzy’ (non-commutative) due to a magnetic monopole at its centre. It offers distinct advantages including the exact preservation of rotation symmetry, the direct observation of emergent conformal symmetry and the efficient extraction of conformal data. In the fuzzy sphere method, the state-operator correspondence plays an essential role. Specifically, there is a one-to-one correspondence between the eigenstates of the critical Hamiltonian on the sphere and the CFT operators, where the energy gaps are proportional to the scaling dimensions. The power of this approach has been first demonstrated in the context of the 3D Ising transition [23], where the presence of emergent conformal symmetry has been convincingly established. Since its proposal, various studies have greatly expanded its horizon [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], including
@@
    accessing the ‘conformal data’ of the 3D Ising CFT such as the OPE coefficients [24, 42], correlation functions [25], entropic $F$-function [31], conformal generators [36, 37] and crosscap coefficients [50],
@@
    studying the conformal defects, boundaries such as the magnetic line defect [27], its $g$-function and defect changing operators [30], its cusp [32], and various conformal boundaries in the 3D Ising CFT [33, 35],
@@
    realising various 3D CFTs such as the Wilson-Fisher CFTs [29], $\mathrm{SO}(5)$ [26] and $\mathrm{O}(4)$ [48] deconfined criticality, a series of new $\mathrm{Sp}(N)$ CFTs [38], Potts model [40], Yang-Lee criticality [43, 44, 45], free real scalar [46, 47] and Chern-Simons-matter CFTs [49].
@@
All the CFTs that have been realised on fuzzy sphere so far have purely bosonic operator content. Apart from the fuzzy sphere regularisation, a recent exploratory attempt to access the fermionic Gross-Neveu-Yukawa CFT considered Dirac fermions on a conventional sphere, using a basis of spinor spherical harmonics [51]. Although intuitive, this approach suffers from UV divergences similar to those in standard QFTs on continuous space, and thus further theoretical and numerical efforts may be required to justify its validity. In contrast, the fuzzy sphere approach, similar to the highly successful lattice regularisation of QFTs and strongly interacting condensed matter lattice models, is free of UV divergences [46]. [^1] In the continuum limit (i. e. the monopole flux $q\to\infty$) of the fuzzy sphere model, any (bare) local operator, as well as its correlators, remain non-diverging. Moreover, the microscopic couplings at the critical point also remain finite, i. e., they converge to finite values in the continuum limit. Consequently, the study of fermionic CFTs would greatly benefit from a fuzzy sphere realisation, especially in light of the success achieved for their bosonic counterparts.
@@
Nevertheless, the fuzzy sphere realisation of fermionic CFTs faces some challenges. Most fuzzy sphere models are built of microscopic fermions, i. e. the fermions with $\mathrm{U}(1)$ electric charge on the lowest Landau level. These microscopic fermions experience the magnetic field generated by the monopole and thus become fuzzy, in the sense that they have strong non-commutativity inherited from the magnetic field. For example, in flat space the microscopic fermions transform in a projective representation of the continuous translation group, which is incompatible with any local operator of a CFT. The resolution is that these microscopic fermions must be gapped, so that the electric charge $\mathrm{U}(1)$ is decoupled from the IR critical theory. Gapless ‘local’ CFT operators on the fuzzy sphere tend to involve bilinear expressions in the microscopic fermions, which are charge neutral and bosonic. Moreover, another challenge is that the fermionic field in the CFT should carry half-integer spin (angular momentum), a requirement that microscopic fermions may not automatically satisfy. In this paper, we show that these challenges can be overcome by a fuzzy sphere model with boson-fermion mixture. Concretely, we consider a setup with both microscopic bosons $b$ and fermions $f$ with the same electric charge and let their angular momenta differ by $1/2$. Their bilinears $b^{\dagger}f$ or $f^{\dagger}b$ are neutral under electric charge and have fermionic statistics as well as half-integer spins. Therefore, the bilinears can be used to realise fermionic local operators in a CFTs. [^2] A recent work [49] have also studied a boson-fermion mixture fuzzy sphere model that realises the confinement transition of $\nu=1/2$ bosonic Laughlin state. Its setup involves two flavours of charge-1 microscopic fermions and a flavour of charge-2 microscopic bosons. Hence, the charge-0 CFT operators are still bosonic.
@@
The simplest fermionic CFT to target is the free Majorana fermion CFT. Its Lagrangian is
@@
where $\chi$ is a two component spinor, $\bar{\chi}=\chi^{\mathrm{T}}\epsilon$, $\epsilon=i\sigma^{2}$, and $\not{\partial}=\sigma^{\mu}\partial_{\mu}$. The free Majorana fermion $\chi$ has exact scaling dimension $\Delta_{\chi}=1$ and is subject to the equation of motion $\not{\partial}\chi=0$. [^3] Unitarity and conformal symmetry imply that any local operator of spin-$\tfrac{1}{2}$ with $\Delta=1$ is a free fermion, so this scaling dimension also gives a clear criterion to identify the free fermion CFT. It describes a transition driven by the mass $m$. Masses with different signs flow respectively to a trivially gapped phase and a $p$-wave chiral topological superconductor [52, 53] (TSC), which is a gapped phase characterised by a parity and time-reversal breaking fermion pairing, thermal Hall conductance or chiral central charge $c_{-}=1/2$ and a chiral free Majorana fermion mode on the edge. [^4] Albeit the name ‘superconductor,’ the TSC do not spontaneously break any $\mathrm{U}(1)$ symmetry.
@@
In this paper, we study the setup with one flavour of fermions and one flavour of bosons on the fuzzy sphere with the same electric charge but a mismatch $1/2$ in the angular momenta. We set the total filling $\nu=1$ and allow the conversion between two fermions and two bosons. By tuning a relative chemical potential, we identify a fermionic integer quantum Hall (fIQH) phase at $\nu_{f}=1$ when fermions are energetically favourable, a bosonic Pfaffian (bPf) phase [52, 54] at $\nu_{b}=1$ with topological order characterised by a $\mathrm{SU}(2)_{2}$ Chern-Simons theory when bosons are energetically favourable, and an intermediate gapped phase corresponding to an $f$-wave TSC (i. e. a $p$-wave TSC on the background of a $\nu_{f}=1$ fIQH state). We show theoretically and numerically that the fIQH-TSC transition is described by a free Majorana fermion CFT. In particular, the energy gap, operator spectrum and two-point correlation function of local fermion all agree well with the free Majorana fermion CFT.
@@
On the other hand, the TSC-bPf transition is described by an Ising CFT coupled to a $\mathbb{Z}_{2}$ gauge field. The gauged Ising universality, also known as Ising, was first studied in the confinement transition of $\mathbb{Z}_{2}$ topological order [55, 56, 57, 58, 59, 60, 61]. While the Ising criticality was first proposed to describe the transition between a trivially gapped state and a $\mathbb{Z}_{2}$ topological order, in our setup, confining a $\mathbb{Z}_{2}$ gauge field drives the bPf phase to an TSC with trivial topological order. The only difference is a TSC background that does not change the universality. This is numerically confirmed by the agreement of the operators spectrum with the $\mathbb{Z}_{2}$-even sector of the Ising CFT. An interesting feature of the gauged Ising CFT is the topological Wilson line defect. We show that a $\mathbb{Z}_{2}$ gauge charge as the endpoint of the Wilson line can be inserted by inserting a $2\pi$ monopole flux, and the endpoint operators correspond to the $\mathbb{Z}_{2}$-odd sector of the Ising CFT.
@@
In short, our boson-fermion mixture construction successfully extends the fuzzy sphere approach to fermionic CFTs, opening avenues to explore a broad class of theories, including the Gross-Neveu-Yukawa, super-Ising, and quantum Hall transitions relevant to Moiré materials. A key ingredient is to create a half-integer fermionic field through the angular momentum offset of $1/2$ between microscopic fermions and bosons. This idea might be generalisable to define CFTs such that the most relevant operators have non-trivial spin: a ‘pure symmetry’ CFT such that the simplest local operators are conserved currents of spin-$1$, or a ‘pure energy’ CFT such that the simplest local operator is the stress tensor, a property expected from holographic duals to theories of quantum gravity [62, 63, 64]. Both would be examples of ‘self-organised’ critical theories stable to any deformation. It may also give a route to engineer spin-1 gauge fields and thus implement general renormalisable Lagrangians of quantum field theories on the fuzzy sphere. It may also be possible to reach effective field theories with higher spin quantum fields. The realisation of the gauged Ising CFT provides new possibilities in studying line defects like monodromy defect in Ising CFT. Last but not least, we remark that the entire phase diagram of our model, including the three gapped phases and two critical points, can be captured by a single Lagrangian of adjoint QCD [65]. A natural future direction is to explore whether a similar setup can be used to study other gauge theories that exhibit rich phase diagrams as well as novel interacting CFTs.
@@
This paper is organised as the following: In Section 2, we present the setup of fuzzy sphere and our model Hamiltonian. In Section 3, we present the phase diagram supported by the numerical evidences for the gapped phases. In Section 4, we analyse the critical theories that describe the two continuous transitions between the three gapped phases. Then, Sections 5 and 6 are devoted to the numerical results respectively for the free Majorana fermion CFT and the gauged Ising CFT. Finally, in Section 7, we make a brief summary and discuss the implications of our work and future directions.

### v3

# Free and Interacting Fermionic Conformal Field Theories on the Fuzzy Sphere

@@ Abstract
The fuzzy-sphere regularisation is a powerful tool to study conformal field theories (CFT) in three spacetime dimensions. In this paper, we extend its scope to CFTs with local fermionic operators. We realise the free-Majorana-fermion CFT on a set-up with one flavour of bosons and one flavour of fermions on the lowest Landau level with a $1/2$ angular momentum mismatch and allow conversion between two bosons and two fermions, and use a relative chemical potential as the tuning parameter. On the phase diagram, we observe two continuous transitions described respectively by a free Majorana fermion and a gauged Ising CFT. We numerically confirm the emergent conformal symmetry through the operator spectrum and the two-point correlation function of the local Majorana fermion. We further establish a correspondence between the fuzzy-sphere models and the field-theory Lagrangians, and extend it to an interacting fermionic CFT — the super-Ising theory with emergent super-conformal symmetry.
@@ 1 Introduction
Conformal field theory (CFT) is one of the central topics of modern physics. CFT has provided important insights into various aspects of theoretical physics. In condensed-matter physics, it has produced useful predictions about the critical phenomena [1, 2, 3]. Many classical and quantum phase transitions are conjectured to have emergent conformal symmetry in the infra-red (IR). In high-energy physics and quantum field theories, CFT also plays important role in understanding string theory [4], duality with gravitational theories on anti-de Sitter (AdS) space [5], and renormalisation group flow structure [6]. In 2D, the infinite-dimensional conformal algebra has made many theories exactly solvable [7, 8, 9]. However, going to the higher dimensions, the CFTs are less well-studied due to a much smaller conformal group, although some of the conformal data are determined at high precision by conformal bootstrap [10, 11].
@@
Fermionic CFTs, i. e. CFTs that include fermionic local operators, constitute a particularly interesting type of CFTs, including Gross-Neveu-Yukawa theories and certain gauge theories with or without Chern-Simons (CS) term. In some special cases, these CFTs can have emergent supersymmetry. For example, the theory of a single Majorana fermion $\chi$ coupled to a real Yukawa scalar field $\sigma$ flows to a super-conformal field theory (SCFT) with $\mathscr{N}=1$ in the IR, known as the super-Ising theory [12, 13, 14, 15, 16], where $\sigma$, $\chi$ and $\epsilon=\sigma^{2}$ combine into a super-conformal multiple and their scaling dimensions are locked together. Experimentally, these fermionic CFTs are easier to detect than their bosonic counterparts, as they produce rich electric signals, to which most experimental probes are naturally sensitive. For example, the frequency dependence of conductivity can be measured through electric transport experiments, and it corresponds to the OPE of the conserved symmetry current $J\times J$ [17, 18], including the central charge $C_{J}$ from its two-point function. An outstanding platform is the Moiré materials [19, 20, 21]. Through tuning many available knobs like potential amplitude, they can be used to realise plateau transitions between fractional Chern insulators [22, 23, 24]. However, the nature and critical data of these theories remain poorly understood.
@@
The fuzzy-sphere regularisation [25] has emerged as a new powerful method to study 3D CFTs. This method involves studying quantum systems on a sphere that is ‘fuzzy’ (non-commutative) due to a magnetic monopole at its centre [26]. It offers distinct advantages including the exact preservation of rotation symmetry, the direct observation of emergent conformal symmetry and the efficient extraction of conformal data. In the fuzzy-sphere method, the state-operator correspondence plays an essential role. Specifically, there is a one-to-one correspondence between the eigenstates of the critical Hamiltonian on the sphere and the CFT operators, where the energy gaps are proportional to the scaling dimensions. The power of this approach has been first demonstrated in the context of the 3D Ising transition [25], where the presence of emergent conformal symmetry has been convincingly established. Since its proposal, various studies have greatly expanded its horizon, including
@@
    accessing the ‘conformal data’ of the 3D Ising CFT, such as the OPE coefficients [27], correlation functions [28], entropic $F$-function [29], conformal generators [30, 31] and cross-cap coefficients [32],
@@
    developing and applying techniques to improve the numerical precision, such as quantum Monte Carlo simulation [33] and conformal perturbation [34],
@@
    realising various 3D CFTs at integer filling, such as the free scalar [35, 36] and Wilson-Fisher [37] CFTs, critical gauge theories realised through non-linear $\sigma$ model including $\mathrm{SO}(5)$ [38, 39] and $\mathrm{O}(4)$ [40] deconfined criticality and a series of new $\mathrm{Sp}(N)$ CFTs [41], Potts model [42], and Yang-Lee criticality [43, 44, 45],
  4. 4.
    exploring fractional quantum Hall (FQH) transitions, such as the Ising CFT on a FQH background [46] and the confinement transition of the $\nu=1/2$ bosonic Laughlin state [47], and
  5. 5.
    studying the conformal defects (e. g. the pinning-field defect [48] and its $g$-function [49] and cusp [50]), conformal boundaries [51, 52] and lower-dimensional CFTs [53].
@@
All the CFTs that have been realised on fuzzy sphere so far have purely bosonic operator content. Apart from the fuzzy-sphere regularisation, a recent exploratory attempt to access the fermionic Gross-Neveu-Yukawa CFT considered Dirac fermions on a conventional sphere, using a basis of spinor spherical harmonics [54]. Although intuitive, this approach suffers from ultra-violet (UV) divergences similar to those in standard QFTs on continuous space, and thus further theoretical and numerical efforts may be required to justify its validity. In contrast, the fuzzy-sphere approach, similar to the highly successful lattice regularisation of QFTs and strongly interacting condensed-matter lattice models, is free of UV divergences [35]. [^1] In the continuum limit (i. e. the monopole flux $q\to\infty$) of the fuzzy-sphere model, any (bare) local operator, as well as its correlators, remain non-diverging. Moreover, the microscopic couplings at the critical point also remain finite, i. e., they converge to finite values in the continuum limit. Consequently, the study of fermionic CFTs would greatly benefit from a fuzzy-sphere realisation, especially in light of the success achieved for their bosonic counterparts.
@@
Nevertheless, the fuzzy-sphere realisation of fermionic CFTs faces some challenges. Most fuzzy-sphere models are built of microscopic fermions, i. e. the fermions with $\mathrm{U}(1)$ electric charge on the lowest Landau level. These microscopic fermions experience the magnetic field generated by the monopole and thus become fuzzy, in the sense that they have strong non-commutativity inherited from the magnetic field. For example, in flat space the microscopic fermions transform in a projective representation of the continuous translation group, which is incompatible with any local operator of a CFT. The resolution is that these microscopic fermions must be gapped, so that the electric charge $\mathrm{U}(1)$ is decoupled from the IR critical theory. Gapless ‘local’ CFT operators on the fuzzy sphere tend to involve bilinear expressions in the microscopic fermions, which are charge neutral and bosonic. Moreover, another challenge is that the fermionic field in the CFT should carry half-integer spin (angular momentum), a requirement that microscopic fermions may not automatically satisfy. In this paper, we show that these challenges can be overcome by a fuzzy-sphere model with boson-fermion mixture. Concretely, we consider a set-up with both microscopic bosons $b$ and fermions $f$ with the same electric charge and let their angular momenta differ by $1/2$. Their bilinears $b^{\dagger}f$ or $f^{\dagger}b$ are neutral under electric charge and have fermionic statistics as well as half-integer spins. Therefore, the bilinears can be used to realise fermionic local operators in a CFTs. [^2] A recent work [47] have also studied a boson-fermion mixture fuzzy-sphere model that realises the confinement transition of $\nu=1/2$ bosonic Laughlin state. Its set-up involves two flavours of charge-1 microscopic fermions and a flavour of charge-2 microscopic bosons. Hence, the charge-0 CFT operators are still bosonic.
@@
The simplest fermionic CFT to target is a massless free Majorana fermion. Its Lagrangian is
@@
where $\chi$ is a two component spinor, $\bar{\chi}=\chi^{\mathrm{T}}\epsilon$, $\epsilon=i\gamma^{2}$, [^3] In 3D Euclidean signature, the $\gamma$-matrices are identical to the Pauli matrices. We use $\gamma$ to avoid overloading the symbol $\sigma$. and $\not{\partial}=\gamma^{\mu}\partial_{\mu}$. The free Majorana fermion $\chi$ has exact scaling dimension $\Delta_{\chi}=1$ and is subject to the equation of motion $\not{\partial}\chi=0$. [^4] Unitarity and conformal symmetry imply that any local operator of spin-$\tfrac{1}{2}$ with $\Delta=1$ is a free fermion, so this scaling dimension also gives a clear criterion to identify the free-fermion CFT. It describes a transition driven by the mass $m$. Masses with different signs flow respectively to a trivially gapped phase and a $\nu_{K}=1$ Majorana quantum Hall (MQH) phase [55, 56]. The MQH is a gapped phase with no topological order characterised by thermal Hall conductance or chiral central charge $c_{-}=1/2$ and a chiral free Majorana fermion mode on the edge. The Kitaev index $\nu_{K}$ appears in the ‘16-fold way’ classification of topological superconductors from spinless fermion [57] and gives the number of chiral edge Majorana fermions and the chiral central charge $c_{-}=\nu_{K}/2$. [^5] In this language, the $\nu=1$ fermionic integer quantum Hall phase can be regarded as a MQH phase with $\nu_{K}=2$. It is also known as the $p$-wave topological chiral superconductor (TSC). [^6] The name captures the parity and time-reversal breaking fermion pairing, although the TSC do not spontaneously break any $\mathrm{U}(1)$ symmetry.
@@
In this paper, we study the set-up with one flavour of fermions and one flavour of bosons on the fuzzy sphere with the same electric charge but a mismatch $1/2$ in the angular momenta. We set the total filling $\nu=1$ and allow the conversion between two fermions and two bosons. By tuning a relative chemical potential, we identify a fermionic integer quantum Hall (fIQH) phase at $\nu=1$ when fermions are energetically favourable, a bosonic Pfaffian (bPf) phase [55, 58] at $\nu_{b}=1$ with topological order characterised by a $\mathrm{SU}(2)_{2}$ Chern-Simons theory when bosons are energetically favourable, and an intermediate gapped phase corresponding to a $\nu_{K}=3$ MQH (i. e. a $\nu_{K}=1$ MQH on the background of a $\nu=1$ fIQH state, or a $f$-wave TSC). We show theoretically and numerically that the fIQH-MQH transition is described by a free Majorana fermion. In particular, the energy gap, operator spectrum and two-point correlation function of local fermion all agree well with the free-Majorana-fermion CFT.
@@
On the other hand, the MQH-bPf transition is described by an Ising CFT coupled to a $\mathbb{Z}_{2}$ gauge field. The gauged Ising universality, also known as Ising, was first studied in the confinement transition of $\mathbb{Z}_{2}$ topological order [59, 60, 61, 62, 63, 64, 65]. While the Ising criticality was first proposed to describe the transition between a trivially gapped state and a $\mathbb{Z}_{2}$ topological order, in our set-up, confining a $\mathbb{Z}_{2}$ gauge field drives the bPf phase to a MQH with trivial topological order. The only difference is a MQH background that does not change the universality. This is numerically confirmed by the agreement of the operators spectrum with the $\mathbb{Z}_{2}$-even sector of the Ising CFT. An interesting feature of the gauged Ising CFT is the topological Wilson line defect. We show that a $\mathbb{Z}_{2}$ gauge charge as the end-point of the Wilson line can be inserted by inserting a $2\pi$ monopole flux, and the end-point operators correspond to the $\mathbb{Z}_{2}$-odd sector of the Ising CFT.
We observe an analogy between the fuzzy-sphere models and the QFT Lagrangians in the realisation of both free Majorana fermion and critical real scalar. In particular, a flavour serves as the reference state from which the CFT ground state can be regarded as a perturabtion. The elementary fields that constitute the field theory are identified as hopping from the reference flavour to the excitation flavours. The interaction terms in the Hamiltonian can be identified with the terms in the QFT Lagrangian. We extend this correspondence to a simple interacting fermionic theory — the super-Ising theory with a Majorana fermion coupled to a real scalar through a Yukawa interaction. We provide numerical evidence for the emergence of super-conformal symmetry and the agreement with the conformal bootstrap results.
@@
In short, our boson-fermion mixture construction successfully extends the fuzzy-sphere approach to free and interacting fermionic CFTs, opening avenues to explore a broad class of theories, including the Gross-Neveu-Yukawa and quantum Hall transitions relevant to Moiré materials. A key ingredient is to create a half-integer fermionic field through the angular momentum offset of $1/2$ between microscopic fermions and bosons. This idea might be generalisable to define CFTs such that the most relevant operators have non-trivial spin: a ‘pure-symmetry’ CFT such that the simplest local operators are conserved currents of spin-$1$, or a ‘pure-energy’ CFT such that the simplest local operator is the stress tensor, a property expected from holographic duals to theories of quantum gravity [66, 67, 68]. Both would be examples of ‘self-organised’ critical theories stable to any deformation. It may also give a route to engineer spin-1 gauge fields and thus implement general renormalisable Lagrangians of quantum field theories on the fuzzy sphere. It may also be possible to reach effective field theories with higher spin quantum fields. The realisation of the gauged Ising CFT provides new possibilities in studying line defects like monodromy defect in Ising CFT.
@@
This paper is organised as the following: In Section 2, we present the set-up of fuzzy sphere and our model Hamiltonian. In Section 3, we present the phase diagram supported by the numerical evidences for the gapped phases. In Section 4, we analyse the critical theories that describe the two continuous transitions between the three gapped phases. Then, Sections 5 and 6 are devoted to the numerical results respectively for the free-Majorana-fermion CFT and the gauged Ising CFT. In Section 7, we establish the correspondence between fuzzy-sphere models and QFT Lagrangians and discuss the realisation of the super-Ising theory in light of this, and present in Section 8 the numerical evidence for the emergent super-conformal symmetry. Finally, in Section 9, we make a brief summary and discuss the implications of our work and future directions.